# Forge: Foundational Optimization Representations from Graph Embeddings

### v2

@@ Abstract
Combinatorial optimization problems are ubiquitous in science and engineering, yet learning-based approaches to accelerate their solution often require solving a large number of hard-to-solve optimization instances to collect training data, incurring significant computational overhead. Existing methods require training dedicated models for each problem distribution for each downstream task, severely limiting their scalability and generalization. In this work, we introduce Forge, a method of pre-training a vector-quantized graph autoencoder on a large and diverse collection of mixed-integer programming (MIP) instances in an unsupervised fashion without dependency on their solution. The vector quantization process creates discrete code assignments that act as a vocabulary to represent optimization instances. We evaluate our approach under both supervised and unsupervised settings. For the unsupervised setting, we demonstrate that Forge embeddings effectively differentiate and cluster unseen instances. For the supervised setting, we fine-tune Forge embeddings and show that a single model predicts both the variables for warm-starts and integrality gaps for cut-generation across multiple problem type distributions. Both predictions help improve performance of a state-of-the-art, commercial optimization solver. Finally, we release our code and pre-trained Forge weights to encourage further research and practical use of instance-level MIP embeddings at https://github.com/skadio/forge/.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are fundamental in science and engineering with widespread applications across multiple domains, including logistics, energy systems, network design, and recommendations (Gabriel Crainic et al., 2021; Miller et al., 2023; Kadıoğlu et al., 2024). Traditionally, CO problems have been addressed using carefully designed meta-heuristics and sophisticated optimization solvers. These classical approaches often demand significant domain expertise and computational resources, especially when problem size and complexity scales up.
@@
Recent advances in Machine Learning (ML) have introduced promising alternatives for solving CO problems. The approaches for ML-guided optimization fall mainly under two categories: 1) end-to-end models that attempt to predict solutions (or objective) without depending on meta-heuristics or solvers, and 2) hybrid methods that 2a) replace computationally intensive components of traditional optimization solvers with learned models, e.g., learning to predict strong branching heuristics (Gasse et al., 2019), and 2b) guide meta-heuristics with learning from feedback (Cai et al., 2025b). For a comprehensive overview of ML-guided CO, we refer readers to the survey by Bengio et al. (2021).
@@
Despite their potential, many of these learning-based methods face practical limitations. A significant drawback of these learning approaches is their heavy dependency on offline training. Training is computationally costly, depends on carefully curating training datasets with desired properties and distributions of the underlying CO instances, and has limited generalization. Moreover, training often depends on using optimization solvers in the first place to create labeled datasets, which defeats the purpose of improving solving for very hard instances that optimization solvers cannot deal with today. Adapting learning-based methods to new distributions and domains remains a challenge, and hence, designing foundational optimization representations in an unsupervised fashion applicable to multiple optimization tasks are a much-needed alternative. This is what we study in this paper.
@@
Our main motivation is grounded in the exceedingly successful ML methods in other modalities such as foundational text and image embeddings. This raises a natural question for optimization: can we leverage the large number of publicly available mixed-integer programming (MIP) instances to develop a pre-trained, general-purpose, foundational model for MIP representation that is useful across multiple optimization tasks across varying sizes and problem domains?
@@
The success of current ML-based approaches for optimization problems makes this direction imminent. For example, Zhou et al. (2023) proposes a meta-learning framework capable of generalizing across variants of vehicle routing problems of different sizes, albeit it is limited to one specific optimization problem, i.e., routing. Similarly, Cai et al. (2025a) introduces a multi-task framework based on a shared model for backdoor prediction and solver configuration, however, the shared model is trained separately per problem. Likewise, Li et al. (2025) proposes an LLM-based evolutionary framework to generate a large set of diverse MIP problems, but remains supervised, and requires a large number of apriori solved instances. While these are promising, the existing work either generalizes across multiple tasks but remains specific to one problem type, or scales to different sizes and problem variants but is specific to one task, or remains supervised. We envision a foundational model to create MIP embeddings that are generalizable to multiple tasks, across different problems, trained in an unsupervised fashion, without dependency on solving hard optimization problems.
@@
When we consider Natural Language Processing (NLP) and Computer Vision (CV), foundational models have emerged through unsupervised or self-supervised training due to the abundance of data. CO problems, while also having a large amount of data, span highly heterogeneous problem types (e.g., Set Cover versus Combinatorial Auction) and present significant variability within each problem type. Most ML-based approaches to CO rely on Graph Neural Networks (GNN), as proposed in (Gasse et al., 2019), which are effective at capturing local variable- and constraint-level information but struggle to capture meaningful global information at the instance level, as studied in detail by Feng et al. (2025) where it was shown that GNNs often struggle to capture the global structure necessary for solving CO problems due to their inherent locality bias. As such, while embeddings for several other modalities such as words, images, and audio exist and have become widespread, to date, there exists no general-purpose MIP embeddings at the instance-level.
@@
With this vision in mind, we propose Forge: Foundational Optimization Representations from Graph Embeddings. Forge is a foundational model to generate useful MIP embeddings through a pre-training framework that learns structural representations at the instance level from a broad distribution of MIP instances without requiring access to their solutions.
@@
To achieve this, we incorporate two key ideas; one originating from NLP and the other from CV. As in NLP, we seek to build a vocabulary to represent the latent space of optimization problems, which enables instance-level representations. As in CV, we leverage vector quantization to preserve global information, which overcomes the issue of GNNs used for CO in previous works. By extending these two crucial insights into the CO context, we make the following concrete contributions:
@@
  • We propose Forge, a general-purpose foundational model to generate MIP embeddings. We show that Forge effectively captures both local and global structures critical to optimization. Unlike previous work, with a single pre-trained Forge model, we provide embeddings at various levels; both instance-level representations, i.e., a single vector representation per MIP instance, and fine-grained variable- and constraint-embeddings.
@@
  • For the unsupervised setting, we show that our instance level MIP embeddings clusters previously unseen instances across multiple problem types with high accuracy.
  • For the supervised setting, we show that pre-trained Forge embeddings can be fine-tuned for various downstream tasks using a small number of labeled data. We evaluate Forge on two radically different tasks to showcase its versatility: predicting variables for warm-starts and predicting integrality gap for cut-generation. Notably, a single pre-trained Forge model is fine-tuned and applied across diverse problem types including Combinatorial Auctions (CA), Set Covering (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC).
@@
  • To enhance traditional optimization solvers, we integrate predictions from Forge into Gurobi (Gurobi Optimization, LLC, 2024), a state-of-the-art commercial solver, and demonstrate consistently lower primal gaps across a range of problem domains and sizes.
  • To enhance ML-guided optimization methods, we incorporate Forge embeddings into PS-Gurobi (Han et al., 2023), enhancing its performance.
The remainder of the paper is structured as follows: we start with a brief background of CO in Section §2 followed by the Forge methodology in Section §3. Sections §4 and §5 show experiments to demonstrate the efficacy of Forge in unsupervised and supervised settings respectively.
@@
Figure 1: Forge: The overall unsupervised approach for learning MIP embeddings. Forge uses a vector quantized graph autoencoder (C-D) to reconstruct node features and edges and is pre-trained across a diverse set of problem types and difficulty levels, enabling it to learn generalizable structural representations without requiring access to optimal solutions.
@@ 2 Background: Mixed-Integer Programming
We deal with solving Mixed-Integer Programming (MIP) instances that formulates combinatorial optimization problems of the form:
$f(x)=\min\{c^{T}x\mid Ax\leq b,x\in\mathbb{R}^{n},x_{j}\in\mathbb{Z}\ \forall j\in I\}$
where $f(x)$ is the objective function, and $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m},c\in\mathbb{R}^{n}$, and the non-empty set $I\subseteq{1,...,n}$ indexes the integer variables.
The Linear Programming (LP) relaxation of a MIP $x_{lp}$ is obtained by relaxing integer variables to continuous variables, i.e., by replacing the integer constraint $x_{j}\in\mathbb{Z}\ \forall j\in I$ to $x_{j}\in\mathbb{R}\ \forall j\in I$. The LP relaxation is an essential part of the branch-and-bound algorithm for providing bounds.
The integrality gap measures how much worse the optimal solution of the LP relaxation when compared to the optimal solution of the original MIP.
3 Forge: Unsupervised Approach to Learn MIP Embeddings
MIP instances are typically represented as a bipartite graph between variables and constraints augmented with node features (Gasse et al., 2019; Ferber et al., 2022; Yau et al., 2024; Chen et al., 2023). This is then followed by training a GNN in a supervised fashion for a specific downstream task on a certain problem class. For example, in Han et al. (2023), the GNN is used for predicting variables for warm-starts trained on Set Cover (SC) and Independent Set (IS) problems. Similarly, in Li et al. (2025), a GNN is used for predicting the integrality gap. Numerous variants follow this template (see §6). Notice that, all of these methods require supervision and do not yield a general-purpose MIP embeddings at the instance level. Taking this a step further; our goal is to learn the structure of MIP instances in an unsupervised manner.
@@ 3 Forge: Unsupervised Approach to Learn MIP Embeddings
A) MIP-to-BP: Given a MIP instance, we start with its bipartite (BP) representation and node features. Each node in this bipartite graph represents a constraint or a variable, with edges indicating which variables are part of which constraints. Each node is also associated with node features and each edge is weighted by the coefficient of the variable in the constraint. Node features are typically extracted from the internal branch-and-bound search tree when solving an instance. For example, there are 18 input node features used in Gasse et al. (2019). We do not attempt to solve or depend on the solution of the instance. Instead, Forge only uses the basic properties of the input instance. For each constraint node, we introduce 4 features composed of its sense (i.e., >>, << or ==) and the RHS value. For each variable node, we introduce 6 features composed of its type (integer, binary, continuous), upper/lower bound, and the coefficient in the objective function. In total, for each node, we obtain a vector of size 10, padded with zeros accordingly based on node type (Figure 1(A)).
B) BP-to-GNN: This bipartite graph with 10-dimensional input node features is then passed into a GNN, similar to previous works, to generate embeddings for each constraint and variable node. More specifically, Forge uses two GraphSage (Hamilton et al., 2017) layers that project each input node into a $d$ dimensional embedding space.
@@
As discussed earlier, while GNNs are good at capturing local variable- and constraint-level information, they struggle to capture meaningful global information at the instance level due to their inherent locality bias (Feng et al., 2025). Preserving global structure is important in CO problems, especially to generalize across problem types (Figure 1(B)).
@@
C) Vector Quantized Codebook: To preserve global structure, we introduce a vector quantized codebook with $k$ discrete codes. These codes act as a “vocabulary”, akin to language models, across MIP instances of various domains and difficulties, thereby preserving global structure. The design follows the approaches developed in computer vision (Van Den Oord et al., 2017; Yu et al., 2022; Lee et al., 2022) and the structure-aware graph tokenizer extension in Yang et al. (2024) (Figure 1(C)).
@@
D) GNN-to-CW: GNN embeddings are passed into a vector quantizer which consists of a codebook with $k$ codes. The codebook maps each variable and constraint node to a discrete code. Each code is then mapped into a $d$ dimensional codeword (CW), yielding constraint and variable CW representations. These representations are the same dimension as the hidden GNN layers (Figure 1(D)).

### v4

@@ Abstract
Combinatorial optimization problems are ubiquitous in science and engineering. Still, learning-based approaches to accelerate combinatorial optimization often require solving a large number of difficult instances to collect training data, incurring significant computational cost. Existing learning-based methods require training dedicated models for each problem distribution, for each downstream task, severely limiting their scalability and generalization. We introduce Forge: Foundational Optimization Representations from Graph Embeddings, a framework that pre-trains a vector-quantized graph autoencoder on a large, diverse collection of mixed-integer programming (MIP) instances in an unsupervised manner, without relying on optimization solvers or optimal solutions. Vector quantization produces discrete code assignments that serve as a vocabulary for representing optimization instances. We evaluate Forge in both unsupervised and supervised settings. In the unsupervised setting, Forge embeddings effectively cluster unseen instances across problem domains and sizes. In the supervised setting, we fine-tune Forge embeddings and show that a single pre-trained model helps predicting both the integrality gap for cut-generation and variable hints for search guidance across multiple problem and size distributions. In both tasks, we improve the performance of a commercial optimization solver and outperform state-of-the-art learning-based methods. Finally, we open-source our training code, pre-trained Forge weights, and embeddings for multiple MIP distributions to foster further research in representation learning for optimization problems.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are fundamental in science and engineering with applications spanning diverse domains, including logistics, energy systems, network design, and recommendations (Gabriel Crainic et al., 2021; Miller et al., 2023; Kadıoğlu et al., 2024). Traditionally, CO problems have been solved using carefully designed meta-heuristics and sophisticated optimization solvers. While effective, such classical methods demand significant domain expertise and computational resources, especially as problem size and complexity grow.
@@
Recent advances in Machine Learning (ML) have introduced promising alternatives for solving CO problems. The approaches for ML-guided optimization fall mainly under two categories: 1) end-to-end models that predict solutions or objective without relying on solvers or meta-heuristics, and 2) hybrid methods which either 2a) replace computationally intensive solver components with learned models, e.g., learning to predict strong branching heuristics (Gasse et al., 2019), or 2b) guide meta-heuristics with learning from feedback (Cai et al., 2025b). For a comprehensive overview of ML-guided CO, we refer readers to the survey by Bengio et al. (2021b).
@@
Despite their potential, learning-based methods face practical limitations. A significant drawback of learning approaches is their heavy dependency on offline training. Training is computationally costly, depends on carefully curating training datasets with desired properties and distributions of the underlying CO instances, and has limited generalization. Ironically, training depends on optimization solvers, that they try to accelerate, to create labeled datasets. This defeats the purpose of improving solving for challenge instances that optimization solvers cannot deal with today. Adapting learning-based methods to new distributions and domains remains a challenge, and hence, designing a foundational optimization representation in an unsupervised fashion applicable to multiple optimization scenarios is a much-needed alternative. This is exactly what we study in this paper.
@@
Our main motivation is grounded in the exceedingly successful foundational methods in other modalities, such as text and image embeddings. This raises a natural question for optimization: can we leverage the abundance of publicly available mixed-integer programming (MIP) instances to develop a pre-trained, general-purpose foundational model for MIP representations that serves multiple optimization tasks, across varying problem domains and sizes?
@@
The growing success of ML-based approaches for optimization problems makes this direction imminent. For example, Zhou et al. (2023) propose a meta-learning framework that generalizes across variants of vehicle routing problems of different sizes, but remains limited to routing. Similarly, Cai et al. (2025a) introduce a multi-task framework for backdoor prediction and solver configuration, however, the shared model is trained separately for each problem. Likewise, Li et al. (2025) propose an LLM-based evolutionary framework to generate diverse MIP problems, but is supervised, and requires a large number of pre-solved instances. While promising, the existing work either generalizes across multiple tasks but remains problem-specific, or scales across different sizes and variants but is task-specific, or remains dependent on solvers. We envision a foundational model that produces MIP embeddings generalizable across tasks, problems, and sizes, trained in an unsupervised fashion, without relying on solving hard optimization problems.
@@
When we look at Natural Language Processing (NLP) and Computer Vision (CV), foundational models have emerged through unsupervised or self-supervised training, enabled by the abundance of data. CO problems, while also benefiting from publicly available datasets, span highly heterogeneous problem types (e.g., Set Covering vs. Combinatorial Auction) and exhibit significant variability within each problem. Most ML-based approaches to CO rely on Graph Neural Networks (GNNs), as proposed in (Gasse et al., 2019), which effectively capture local variable and constraint-level information but struggle to encode meaningful global structure. This limitation, rooted in the inherent locality bias of GNNs, is analyzed in detail by (Feng et al., 2025). As a result, while embeddings for other modalities such as text, image, and audio are now widespread, no general-purpose instance-level embeddings exist for MIPs to date.
@@
With this vision in mind, we propose Forge: Foundational Optimization Representations from Graph Embeddings. Forge is a foundational model designed to generate MIP embeddings through a pre-training framework that learns structural representations at the instance level in an unsupervised manner, using a broad distribution of MIP instances without requiring access to their solutions.
@@
To achieve this, we incorporate two key ideas; one inspired by NLP and the other by CV. From NLP, we adopt the concept of a vocabulary to represent the latent space of optimization problems, enabling instance-level representations. From CV, we leverage vector quantization to preserve global information, addressing the limitations of GNN-based approaches in prior CO work. By extending these two crucial insights into the CO context, we make the following concrete contributions:
@@
  • A Foundational Model for Optimization: We propose Forge, a general-purpose foundational model for generating MIP embeddings (§3). Forge captures both local and global structures critical to optimization. Unlike prior work, a single pre-trained Forge model provides embeddings at multiple levels: instance-level representations (one vector per MIP instance), and fine-grained variable and constraint embeddings.
@@
  • Unsupervised Generalization: In the unsupervised setting, Forge embeddings cluster previously unseen instances across diverse problem types with high accuracy (§4).
  • Supervised Adaptability: In the supervised setting, pre-trained Forge embeddings can be fine-tuned on diverse downstream tasks using minimal additional data and a low-cost labeling strategy that avoids solving to optimality. We evaluate Forge on two distinct tasks: estimating integrality gap for cut generation (5.1) and predicting variables for search guidance (5.2). Notably, a single pre-trained Forge model is fine-tuned and applied across varied domains and problem sizes.
@@
  • Solver Integration: To enhance traditional optimization solvers, we integrate Forge predictions into Gurobi (Gurobi Optimization, LLC, 2024), a state-of-the-art commercial solver, and demonstrate consistently lower primal gap across both tasks and a wide range of problem domains and sizes.
  • ML Augmentation: To enhance ML-guided optimization, we evaluate Forge against (Li et al., 2025) for integrality gap prediction, and PS-Gurobi (Han et al., 2023) for search guidance, improving their performance on large sets of instances they were trained on, yet unseen by Forge.
@@
Figure 1: Forge: Our approach for learning MIP embeddings without supervision. Starting with the bipartite representation and its GNN embedding (A-B), Forge uses a vector quantized graph autoencoder (C-D) to reconstruct node features and edges (E). It is pretrained across a diverse set of problems and sizes to learn generic MIP representations without dependency on optimal solutions.
@@ 2 Background: Mixed-Integer Programming
Let us start with a brief background on Mixed-Integer Programming (MIP) that formulates combinatorial optimization problems of the form: $f(x)=\min\{c^{T}x\mid Ax\leq b,x\in\mathbb{R}^{n},x_{j}\in\mathbb{Z}\ \forall j\in I\}$ where $f(x)$ is the objective function, and $A\in\mathbb{R}^{m\times n},b\in\mathbb{R}^{m},c\in\mathbb{R}^{n}$, and the non-empty set $I\subseteq{1,...,n}$ indexes the integer variables. The Linear Programming (LP) relaxation of a MIP $x_{lp}$ is obtained by relaxing integer variables to continuous variables, i.e., by replacing the integer constraint $x_{j}\in\mathbb{Z}\ \forall j\in I$ to $x_{j}\in\mathbb{R}\ \forall j\in I$. The LP relaxation is an essential part of the branch-and-bound algorithm for providing bounds. The integrality gap measures how much worse the optimal solution of the LP relaxation when compared to the optimal solution of the original MIP.
3 Forge: Unsupervised Representation Learning for MIPs
MIP instances are typically represented as a bipartite graph between variables and constraints augmented with node features (Gasse et al., 2019; Ferber et al., 2022; Yau et al., 2024; Chen et al., 2023). This is then followed by training a GNN in a supervised fashion for a specific downstream task on a certain problem class. For example, in Han et al. (2023), the GNN is used for predicting variables for warm-starts trained on Set Cover (SC) and Independent Set (IS) problems. Similarly, in Li et al. (2025), a GNN is used for predicting the integrality gap. Numerous variants follow this template (see related works in §6 and Appendix A.9). Notice that, all of these methods require supervision and do not yield a general-purpose MIP embeddings at the instance level. Taking this a step further; our goal is to learn the structure of MIP instances in an unsupervised manner.
@@ 3 Forge: Unsupervised Representation Learning for MIPs
A) MIP-to-BP: Given a MIP instance, we start with its bipartite (BP) representation and node features. Each node in this bipartite graph represents a constraint or a variable, with edges indicating which variables are part of which constraints. Each node is associated with node features and each edge is weighted by the coefficient of the variable in the constraint. Node features are typically extracted from the internal branch-and-bound search tree when solving an instance. For example, there are 18 input node features used in (Gasse et al., 2019). We do not attempt to solve or depend on the solution of the instance. Instead, Forge only uses the basic properties of the input instance. For each constraint node, we introduce 4 features composed of its sense (i.e., >>, << or ==) and the RHS value. For each variable node, we introduce 6 features composed of its type (integer, binary, continuous), upper/lower bound, and the coefficient in the objective function. In total, for each node, we obtain a vector of size 10, padded with zeros accordingly based on node type (Figure 1-A).
@@
B) BP-to-GNN: This bipartite graph with 10-dimensional input node features is passed into a GNN, akin to previous works, to generate embeddings for each constraint and variable node. More specifically, Forge uses two GraphSage (Hamilton et al., 2017) layers that project each input node into a $d$ dimensional embedding space. As discussed earlier, while GNNs are good at capturing local variable- and constraint-level information, they struggle to capture meaningful global information at the instance level due to their inherent locality bias (Feng et al., 2025). Preserving global structure is important in CO problems, especially to generalize across problem types (Figure 1-B).
@@
C) Vector Quantized Codebook: To preserve global structure, we introduce a vector quantized codebook with $k$ discrete codes. These codes act as a ‘vocabulary’, akin to language models, across MIP instances of various domains and difficulties, thereby preserving global structure. The design follows the approaches developed in computer vision (Van Den Oord et al., 2017; Yu et al., 2022; Lee et al., 2022) and the structure-aware graph tokenizer extension in Yang et al. (2024) (Figure 1-C).
@@
D) GNN-to-CW: GNN embeddings are passed into a vector quantizer which consists of a codebook with $k$ codes. The codebook maps each variable and constraint node to a discrete code. Each code is then mapped into a $d$ dimensional codeword (CW), producing CW representations for constraints and variables, aligned with the dimensionality of the hidden GNN layers (Figure 1-D).